### v1

@@ 2.1 Original NGPPS Sample and Planet Types
Both Emsenhuber et al. (2021a) and Burn et al. (2021) classify the planets into categories based on their mass at the end of the formation phase. These encompass Earth-like, Super-Earth, Neptunian, Sub-Giant and Giant, detailed in Table 1. Figure 1 shows the distribution of the Monte Carlo variables and number of giant planets found in this system for the and run.
Min. Mass [] |
Max. Mass [] |
Earth |
0.5 |
2 |
Super-Earth |
2 |
10 |
Neptunian |
10 |
30 |
Sub-Giant |
30 |
300 |
Giant |
300 |
@@ 2.2 Correlations between Monte Carlo Variables and Planet Type Abundances
Of primary concern for the study of planet populations throughout the galaxy is the relation between the number of planets in a given category and the chosen parameter, since those expected to vary systematically as a function of galactic environment. For example, Figure 1 indicates that higher values of [Fe/H] and lead to a larger number of giant planets. To get a feeling for the general relations, the correlation coefficients between the planet types and Monte Carlo variables are shown in Figure 2 for the and population. It is easy to see that the more massive planet types are positively correlated with a larger initial disc mass, while low mass planets have a negative correlation. The same holds true for the metallicity. In contrast, correlations with the external photo-evaporation rate and inner disc edge are much weaker for the considered range of values. These results persist for initial values for between 10 and 100, and similarly for lower host star masses, although the correlations become weaker with decreasing host mass.
Full Model |
Reduced Model |
Earth |
4.1 |
3.9 |
2.1 |
2.3 |
Super-Earth |
3.4 |
3.0 |
2.2 |
2.6 |
Neptunian |
0.43 |
0.42 |
0.43 |
0.52 |
Sub-Giant |
0.072 |
0.066 |
0.10 |
0.13 |
Giant |
0.31 |
0.35 |
0.47 |
0.58 |
@@ 2.3 Planet Assignment Model
To populate the star particles in the HESTIA galaxy simulation with planets, we need a model that connects the properties of the host star (or stellar population) with the properties of the planets forming around the star. We construct a simple assignment model that connects the occurrence rate of planets in each category with the Monte Carlo variable values. For this purpose, we determine the number of planets per category (as defined in Table 1) across the 1000 Monte Carlo runs (for a fixed and ). The results are smoothed and interpolated using a -nearest neighbours algorithm with (chosen through 10-fold cross-validation on the original sample).
Considering that the photo-evaporation rate and inner disc radius exert minimal influence on planet classifications, we choose to further reduce the model by incorporating only the initial gas disc mass and [Fe/H] as occurrence rate predictors. This increases interpretability of the model at the cost of slightly reduced predictive power. The predicted occurrence rate of planets for the full model (incorporating all four Monte Carlo variables) and the reduced model (using only and [Fe/H]) are shown in Table 2 for the and population. The reduced model yields slightly fewer low-mass planets and more giant ones, as elevated photo-evaporation rates may shorten disc lifetimes and thus stop gas accretion early. These effects are however well within the variability of the model.
@@ 2.4 Influence of the Number of Embryos and stellar Mass
Of further interest is how the relation of planet populations to the systematically varied quantities. In Figure 3, we show the total number of planet per category in all the simulation runs with a varying number of initial embryos (and fixed host star mass, ). The number of massive planets (Neptunians, Sub-Giants, Giants) are nearly constant for all runs with . The number of Earth-like planets and Super-Earths on the other hand increases significantly with the number of embryos. Figure 4 shows a similar figure for a fixed number of initial embryos () and varying host star mass. Here, the number of Earth-like planets does not depend strongly on the host star mass for masses between 0.1 and 1 . The more massive planets however become more frequent for massive stars. Giant planets in particular only occur for host star masses . This agrees with our previous findings that occurrence rates depend primarily on the available solid mass. Burn et al. (2021) assume that the disc mass scales linearly with the mass of the host star. Low-mass stars therefore are less likely to possess the solid material needed to form giant planets within the Bern model.
@@ 2.5 Comparison with Observations
3 Milky Way–Analogue in HESTIA
HESTIA (High-Resolution Environmental Simulations of The Immediate Area) is a suite of hydrodynamical cosmological simulations designed to replicate the Local Group in its correct cosmographic context (Libeskind et al., 2020). It is based on the Arepo code for gravitational N-body systems and magnetohydrodynamics, while the baryonic physics is handled by the Auriga (Grand et al., 2017) model of galaxy formation.
The initial conditions are derived from cosmographic observations to faithfully reconstruct the Local Group, both in terms of its internal structure and nearby large-scale structures. It accurately recreates prominent features of the local Universe such as the Virgo Cluster and Local Void, including their appropriate masses and relative locations to the Local Group (Tully et al., 2009; Karachentsev & Nasonova, 2010). Centered on the Local Group midpoint, the simulation contains a high-resolution region with a radius of 3-5 Mpc. An ensemble of dark matter-only and low-resolution hydrodynamical simulations were run in order to construct a Local Group-like pair of Milky Way– and Andromeda–Analogues with accurate mass ranges and separations. The three most promising simulations have been re-run at a higher resolution. These analogues closely resemble their observational counterparts in terms of total mass, mass ratio, stellar disc mass, morphology separation, relative velocity, rotation curves, bulge-disc morphology, satellite galaxy stellar mass function, and satellite radial distribution. In terms of their formation history, the MW– and Andromeda–analogues grow slower than than comparable halos in unconstrained (McBride et al., 2009; Libeskind et al., 2020). In the highest-resolution runs, the attained mass resolutions of the individual particles are approximately , , , while the spatial resolution is around pc. It needs to be highlighted that the resolution is not sufficient to resolve individual stars, rather the star particles are treated as single stellar populations with uniform age and metallicity.
In this study, we focus on the Milky Way–Analogue in the highest-resolution 37_11 run, which will be described in more detail in the following sections.
@@ 3.1 General Comparison
As described in the previous section, the Milky Way–Analogues in HESTIA have been purposefully crafted to mirror a galaxy that is Milky Way-like in many properties relevant for galaxy formation and evolution. They were, however, not constructed to be a perfect copy of the Milky Way, and some of the properties that are more relevant for planet formation were understandably not considered in the cosmological context they were built for. We, therefore, compare some of the more local and structural properties of our chosen Milky Way–Analogue in more detail.
The virial mass and stellar masses of the Analogue closely match those of the Milky Way within observational constraints (see Table 3). This is unsurprising, given that the simulations were specifically designed to achieve this result. Assuming a Chabrier (2003) IMF, the stellar mass corresponds to roughly 135 billion stars in the Analogue, of which 2.3 billion stars are main-sequence stars within 5% of the mass of the sun. Libeskind et al. (2020) analyzed the structural parameters of the Analogue by fitting a surface brightness profile, which they modelled as the sum of a Sérsic profile describing the bulge and an exponential profile for the disc. The obtained effective bulge and disc radii are are in good agreements with those of the Milky Way (Table 3). While the Sérsic index of the Milky Way bulge lacks precise constraints, the simulation value of 1.80 is comparable to the observational value of for the Andromeda galaxy (Courteau et al., 2011), the closest comparable spiral galaxy.
Furthermore, as discussed in Section 2.3, a major factor in the planet assignment model is the metallicity of the host star. Within HESTIA, 9 elemental species are tracked throughout the simulation: H, He, C, N, O, Ne, Mg, Si, and Fe. From these, we can calculate the iron abundance [Fe/H] for any individual star particle. We assume all stars within a star particle (a single stellar population) share the same metallicity. The metallicity gradient of the Milky Way–Analogue between 2 and 10 kpc from the galactic center is -0.027 dex/kpc. This is a shallower slope compared to the values determined for the Milky Way in observational studies (Lemasle et al., 2007; Luck & Lambert, 2011; Genovali et al., 2014). Although, a more recent study by Lemasle et al. (2018) found Milky Way values of – dex/kpc and – dex/kpc using F/10 double-mode Cepheids, closer to that of the Analogue. Nevertheless, within the HESTIA Analogue, an average solar neighbourhood metallicity value of [Fe/H] -0.1 to 0 (Haywood, 2001) is reached at a distance of – kpc from the galactic center, in good agreement with the distance of the solar neighbourhood from the Milky Way center ( kpc, McMillan, 2017).
@@ 3.2 Galaxy Components
Within the solar neighbourhood, the stars in the Milky Way are generally categorised into three groups: the thin disc, the thick disc, and the stellar halo (Gilmore & Reid, 1983). These populations are distinguished by their kinematics (Piffl et al., 2014; Sanders & Binney, 2015), stellar age (Schuster et al., 2006), and chemical composition (Bensby et al., 2014; Masseron & Gilmore, 2015; Hawkins et al., 2015). Thin disc stars have lower orbital velocity and higher velocity dispersion compared to stars in the thick disc. Chemically, thick disc stars are metal-poor and -enhanced relative to thin disc stars (Prochaska et al., 2000; Reddy et al., 2003, 2006). Halo stars, in contrast, have even larger velocity dispersion (Bond et al., 2010) and are metal-poor. Near the galactic center, there is also a population of bulge stars characterised by high velocity dispersion (Portail et al., 2015) and a wide range of metallicities, spanning from [Fe/H] -1.5 to [Fe/H] +0.5 (Zoccali et al., 2008). These variations in formation history and chemistry are likely to influence the planet populations within these galactic components (Bashi & Zucker, 2022; Nielsen et al., 2023). Therefore, we decompose the HESTIA Analogue into four categories, broadly consistent with the four Milky Way components.
While observational surveys benefit strongly from the separation of components based on combined kinematics and chemistry criteria (Bashi & Zucker, 2019), we adopt a simpler classification approach in this work. We categorise based on the binding energy and circularity of the star particles, aligning with previous studies on galaxy formation simulations (Abadi et al., 2003; Yu et al., 2023) by employing the the morphological decomposition code MORDOR by Zana et al. (2022). Face- and side-on projection maps of the components are shown in Figure 5.
The determined galaxy components broadly align with anticipated properties and observations of the Milky Way. The bulge, thin disc and thick disc closely match the Milky Way in terms of stellar mass, mean stellar age and mean [Fe/H], as well as minor-to-major axis ratio. Discrepancies are found for the stellar halo, which is more massive and metal-rich in the HESTIA Analogue compared to the Milky Way. This can likely be attributed to a misclassification of some bulge and disc stars based on the purely kinematic decomposition. The key attributes are summarised in Table 4, and some distributions are shown in Figure A.2.
@@ 3.3 Integrating The Planet Model
As discussion in Section 2.5, the assignment model predicts an increase of low-mass planet frequency with stellar mass, which is a behaviour not seen in observations. Since low-mass stars are common than massive ones, this could bias our conclusions if not controlled for. We therefore choose to analyse the planet populations for different host star masses independently, and not across stellar masses. To calculate the number of eligible stars associated with the star particle, we integrate the IMF between , where is the chosen (fixed) host star mass. Since planet formation is modeled only for main-sequence stars, we exclude stars with a lifetime shorter than the particle age.
In short, the planet prescription at fixed host star mass is as follows:
1.
Choose an NGPPS population by fixing and .
2.
Obtain the age, [Fe/H], and mass of the star particle.
3.
Sample the Monte Carlo variable from the distribution given by Emsenhuber et al. (2021b).
4.
Integrate the (normalised) Chabrier (2003) IMF, , in the interval , and multiply by the mass of the star particle to obtain the number of stars.
5.
Remove post-main sequence stars based on the age of the star particle 111We assume a lifetime of 10 Gyr for a star with , and scale the lifetimes of less massive stars by a simple power law with an index of 2.5..
6.
Assign the number of planets per star using the planet assignment model for each category (Section 2.3).
7.
Multiply the number of stars by the number of planets per star.
@@ 3.4 Metal-Poor Stars
The [Fe/H] distribution in the HESTIA Analogue covers a wide range of values, with a majority of particles residing between [Fe/H] = -2.5 and [Fe/H] = 1. This interval is considerably broader than the range considered in the NGPPS database, which is limited to — a range selected to cover the typical metallicities within the solar neighbourhood. In the HESTIA Analogue, 80% of the star particles conform to this range. However, around 48% of the star particles in the thick disc and stellar halo fall beneath the NGPPS lower threshold.
Drawing from trends highlighted in the planet assignment model discussed in Section 2.3, considerably larger disc masses would be needed at such low metallicities to provide sufficient solid material for planet formation, especially for planets more massive than . In addition, observations suggest that protoplanetary discs disperse faster in low-metallicity environments (Yasui et al., 2009; Yasui, 2021).
From an observational standpoint, only 26 planets have been confirmed orbiting stars with [Fe/H] ¡ -0.6 at the time of writing, with a minimum metallicity of [Fe/H] = -1.0, providing very limited information on the effects of low-metallicity environments on planet formation. 222exoplanetarchive.ipac.caltech.edu
In our analysis, we will therefore consider two extreme scenarios. In one, we presume the planet frequencies to be unchanging at extremely low metallicities, adopting the same values as in the [Fe/H] = -0.6 case. In the second scenario, we assume that short disc lifetimes and a lack of solid material render planet formation highly unlikely to impossible, and assign an occurrence rate of zero to those particles.

### v2

@@ 2.1 Original NGPPS sample and planet types
Both Emsenhuber et al. (2021a) and Burn et al. (2021) classify the planets into categories based on their mass at the end of the formation phase. These encompass Earth-like, Super-Earth, Neptunian, Sub-giant, and giant, as detailed in Table 1. Figure 1 shows the distribution of the Monte Carlo variables and number of giant planets found in this system for the and run.
| Min. Mass [] | Max. Mass [] | |
| Earth | 0.5 | 2 |
| Super-Earth | 2 | 10 |
| Neptunian | 10 | 30 |
| Sub-giant | 30 | 300 |
| Giant | 300 |
@@ 2.2 Correlations between Monte Carlo variables and planet type abundances
Of primary concern for the study of planet populations throughout the galaxy is the relation between the number of planets in a given category and the chosen parameter, since those expected to vary systematically as a function of galactic environment. For example, Figure 1 indicates that higher values of [Fe/H] and lead to a larger number of giant planets. To get a feeling for the general relations, the correlation coefficients between the planet types and Monte Carlo variables are shown in Figure 2 for the and population. It is easy to see that the more massive planet types are positively correlated with a larger initial disc mass, while low-mass planets have a negative correlation. The same holds true for the metallicity. In contrast, correlations with the external photo-evaporation rate and inner disc edge are much weaker for the considered range of values. These results persist for initial values for between 10 and 100, and similarly for lower host star masses, although the correlations become weaker with decreasing host mass.
| Full Model | Reduced Model | |
| Earth | 4.1 | 3.9 |
| 6.1 | 6.4 | |
| Super-Earth | 3.4 | 3.0 |
| 6.4 | 7.1 | |
| Neptunian | 0.43 | 0.42 |
| 1.2 | 1.6 | |
| Sub-giant | 0.072 | 0.066 |
| 0.43 | 0.53 | |
| Giant | 0.31 | 0.35 |
| 1.3 | 1.5 |
@@ 2.3 Planet assignment model
To populate the star particles in the HESTIA galaxy simulation with planets, we need a model that connects the properties of the host star (or stellar population) with the properties of the planets forming around the star. We constructed a simple assignment model that connects the occurrence rate of planets in each category with the Monte Carlo variable, where the occurrence rate is defined as:
We note that this quantity differs from the fraction of planet-hosting stars, which compares the total number with systems with a planet of a given type compared to the total number of stars.
We determine the number of planets per category (as defined in Table 1) across the 1000 Monte Carlo runs (for a fixed and ). The results are smoothed and interpolated using a -nearest neighbours algorithm with (chosen via a ten-fold cross-validation on the original sample).
Considering that the photo-evaporation rate and inner disc radius exert minimal influence on planet classifications, we chose to further reduce the model by incorporating only the initial gas disc mass and [Fe/H] as occurrence rate predictors. This increases interpretability of the model at the cost of slightly reduced predictive power. The predicted occurrence rate of planets for the full model (incorporating all four Monte Carlo variables) and the reduced model (using only and [Fe/H]) are shown in Table 2 for the and population. The reduced model yields slightly fewer low-mass planets and more giant ones, as elevated photo-evaporation rates may shorten disc lifetimes and thus stop gas accretion early. These effects are however well within the variability of the model.
@@ 2.4 Influence of the number of embryos and stellar mass
One asepct that merits further interest is how the relation of planet populations to the systematically varied quantities. In Figure 3, we show the total number of planet per category in all the simulation runs with a varying number of initial embryos (and fixed host star mass, ). The number of massive planets (Neptunians, Sub-giants, giants) are nearly constant for all runs with . The number of Earth-like planets and Super-Earths, on the other hand, increases significantly with the number of embryos. Figure 4 shows a similar figure for a fixed number of initial embryos () and varying host star mass. Here, the number of Earth-like planets does not depend strongly on the host star mass for masses between 0.1 and 1 . The more massive planets however become more frequent for massive stars. Giant planets, in particular, only occur for host star masses . This agrees with our previous findings that occurrence rates depend primarily on the available solid mass. Burn et al. (2021) assume that the disc mass scales linearly with the mass of the host star. Low-mass stars therefore are less likely to possess the solid material needed to form giant planets within the Bern model.
@@ 2.5 Comparison with observations
3 Milky Way analogue in HESTIA
High-Resolution Environmental Simulations of The Immediate Area (HESTIA) is a suite of hydrodynamical cosmological simulations designed to replicate the Local Group in its correct cosmographic context (Libeskind et al., 2020). It is based on the Arepo code for gravitational N-body systems and magnetohydrodynamics, while the baryonic physics is handled by the Auriga (Grand et al., 2017) model of galaxy formation.
The initial conditions are derived from cosmographic observations to faithfully reconstruct the Local Group, both in terms of its internal structure and nearby large-scale structures. It accurately recreates prominent features of the local Universe such as the Virgo Cluster and Local Void, including their appropriate masses and relative locations to the Local Group (Tully et al., 2009; Karachentsev & Nasonova, 2010). Centered on the Local Group midpoint, the simulation contains a high-resolution region with a radius of 3-5 Mpc. An ensemble of dark matter-only and low-resolution hydrodynamical simulations were run in order to construct a Local Group-like pair of MW and Andromeda analogues with accurate mass ranges and separations. The three most promising simulations have been re-run at a higher resolution. These analogues closely resemble their observational counterparts in terms of total mass, mass ratio, stellar disc mass, morphology separation, relative velocity, rotation curves, bulge-disc morphology, satellite galaxy stellar mass function, and satellite radial distribution. In terms of their formation history, the MW and Andromeda analogues grow slower than than comparable halos in unconstrained (McBride et al., 2009; Libeskind et al., 2020). In the highest-resolution runs, the attained mass resolutions of the individual particles are approximately , , and , while the spatial resolution is around pc. It ought to be highlighted that the resolution is not sufficient to resolve individual stars, rather the star particles are treated as single stellar populations with uniform age and metallicity.
In this study, we focus on the MW analogue in the highest-resolution 37_11 run, which will be described in more detail in the following sections.
@@ 3.1 General comparison
As described in the previous section, the MW analogues in HESTIA have been purposefully crafted to mirror a galaxy that is MW-like in many properties relevant for galaxy formation and evolution. They were, however, not constructed to be a perfect copy of the MW, and some of the properties that are more relevant for planet formation were understandably not considered in the cosmological context they were built for. We therefore compare some of the more local and structural properties of our chosen MW analogue in more detail.
The virial mass and stellar masses of the analogue closely match those of the MW within observational constraints (see Table 3). This is unsurprising, given that the simulations were specifically designed to achieve this result. Assuming a Chabrier (2003) IMF, the stellar mass corresponds to roughly 135 billion stars in the analogue, of which 2.3 billion stars are main-sequence stars within 5% of the mass of the sun. Libeskind et al. (2020) analyzed the structural parameters of the analogue by fitting a surface brightness profile, which they modelled as the sum of a Sérsic profile describing the bulge and an exponential profile for the disc. The obtained effective bulge and disc radii are are in good agreements with those of the MW (Table 3). While the Sérsic index of the MW bulge lacks precise constraints, the simulation value of 1.80 is comparable to the observational value of for the Andromeda galaxy (Courteau et al., 2011), the closest comparable spiral galaxy.
Furthermore, as discussed in Section 2.3, a major factor in the planet assignment model is the metallicity of the host star. Within HESTIA, 9 elemental species are tracked throughout the simulation: H, He, C, N, O, Ne, Mg, Si, and Fe. From these, we can calculate the iron abundance [Fe/H] for any individual star particle. We assume all stars within a star particle (a single stellar population) share the same metallicity. The metallicity gradient of the MW analogue between 2 and 10 kpc from the galactic centre is -0.027 dex/kpc. This is a shallower slope compared to the values determined for the MW in observational studies (Lemasle et al., 2007; Luck & Lambert, 2011; Genovali et al., 2014). Although, a more recent study by Lemasle et al. (2018) found MW values of – dex/kpc and – dex/kpc using F/10 double-mode Cepheids, closer to that of the analogue. Nevertheless, within the HESTIA analogue, an average solar neighbourhood metallicity value of [Fe/H] -0.1 to 0 (Haywood, 2001) is reached at a distance of – kpc from the galactic centre, in good agreement with the distance of the solar neighbourhood from the MW centre ( kpc, McMillan, 2017).
@@ 3.2 Galaxy components
Within the solar neighbourhood, the stars in the MW are generally categorised into three groups: the thin disc, the thick disc, and the stellar halo (Gilmore & Reid, 1983). These populations are distinguished by their kinematics (Piffl et al., 2014; Sanders & Binney, 2015), stellar age (Schuster et al., 2006), and chemical composition (Bensby et al., 2014; Masseron & Gilmore, 2015; Hawkins et al., 2015). Thin disc stars have lower orbital velocity and higher velocity dispersion compared to stars in the thick disc. Chemically, thick disc stars are metal-poor and -enhanced relative to thin-disc stars (Prochaska et al., 2000; Reddy et al., 2003, 2006). Halo stars, in contrast, have even larger velocity dispersion (Bond et al., 2010) and are metal-poor. Near the galactic centre, there is also a population of bulge stars characterised by high velocity dispersion (Portail et al., 2015) and a wide range of metallicities, spanning from [Fe/H] -1.5 to [Fe/H] +0.5 (Zoccali et al., 2008). These variations in formation history and chemistry are likely to influence the planet populations within these galactic components (Bashi & Zucker, 2022; Nielsen et al., 2023). Therefore, we decompose the HESTIA analogue into four categories, broadly consistent with the four MW components.
While observational surveys benefit strongly from the separation of components based on combined kinematics and chemistry criteria (Bashi & Zucker, 2019), we adopt a simpler classification approach in this work. We categorised objects based on the binding energy and circularity of the star particles, aligning with previous studies on galaxy formation simulations (Abadi et al., 2003; Yu et al., 2023) by employing the morphological decomposition code MORDOR by Zana et al. (2022). Face- and side-on projection maps of the components are shown in Figure 5.
The determined galaxy components broadly align with anticipated properties and observations of the MW. The bulge, thin disc, and thick disc closely match those of the MW in terms of stellar mass, mean stellar age, and mean [Fe/H], as well as the minor-to-major axis ratio. Discrepancies are found for the stellar halo, which is more massive and metal-rich in the HESTIA analogue compared to the MW. This can likely be attributed to a misclassification of some bulge and disc stars based on the purely kinematic decomposition. The key attributes are summarised in Table 4, and some distributions are shown in Figure A.2.
@@ 3.3 Integrating the planet model
As discussion in Section 2.5, the assignment model predicts an increase of low-mass planet frequency with stellar mass, which is a behaviour not seen in observations. Since low-mass stars are common than massive ones, this could bias our conclusions if not controlled for. We therefore chose to analyse the planet populations for different host star masses, , independently, and not across stellar masses. To calculate the number of eligible stars associated with the star particle, we integrate the IMF between , where is the chosen (fixed) host star mass. Since planet formation is modeled only for main-sequence stars, we exclude stars with a lifetime shorter than the particle age.
In short, the planet prescription at fixed host star mass is as follows. First, we choose an NGPPS population by fixing and . Then, we obtain the age, [Fe/H], and mass of the star particle. Next, we sample the Monte Carlo variable from the distribution given by Emsenhuber et al. (2021b). We integrate the (normalised) Chabrier (2003) IMF, , in the interval , and multiply by the mass of the star particle to obtain the number of stars. Post-main sequence stars are removed based on the age of the star particle 111We assume a lifetime of 10 Gyr for a star with , and scale the lifetimes of less massive stars by a simple power law with an index of 2.5.. We then assign the number of planets per star using the planet assignment model for each category (Section 2.3). Finally, we multiply the number of stars by the number of planets per star. 222The code used for the analysis in this work has been compiled into a python module skaro and can be found at github.com/ChrisBoettner/skaro.
@@ 3.4 Metal-poor stars
The [Fe/H] distribution in the HESTIA analogue covers a wide range of values, with a majority of particles residing between [Fe/H] = -2.5 and [Fe/H] = 1. This interval is considerably broader than the range considered in the NGPPS database, which is limited to . This range was selected to cover the typical metallicities within the solar neighbourhood. In the HESTIA analogue, 80% of the star particles conform to this range. However, around 48% of the star particles in the thick disc and stellar halo fall beneath the NGPPS lower threshold.
Drawing from trends highlighted in the planet assignment model discussed in Section 2.3, considerably larger disc masses would be needed at such low metallicities to provide sufficient solid material for planet formation, especially for planets more massive than . In addition, observations suggest that protoplanetary discs disperse faster in low-metallicity environments (Yasui et al., 2009; Yasui, 2021). From an observational standpoint, only 26 planets have been confirmed orbiting stars with [Fe/H] ¡ -0.6 at the time of writing, with a minimum metallicity of [Fe/H] = -1.0, providing very limited information on the effects of low-metallicity environments on planet formation. 333exoplanetarchive.ipac.caltech.edu
Therefore, in our analysis, we consider two extreme scenarios. Firstly, we presume the planet frequencies to be unchanging at extremely low metallicities, adopting the same values as in the [Fe/H] = -0.6 case. Secondly, we assume that short disc lifetimes and a lack of solid material render planet formation highly unlikely to impossible, and assign an occurrence rate of zero to those particles.